# Topological Spatial Graph Coarsening

### v1

@@ 6.3 Application on a classification task
In this section, we evaluate our topological and optimized coarsening procedure on a fungi network dataset. We perform a graph classification task on the original and the coarsed data to evaluate the amount of retained and loss information.
We consider a real-world dataset made from graphs representing fungi, extracted from [fricker2025fungi]. Each fungus is represented as a spatial graph where edges represent mycelia [^1] The mycelium is the root-like structure of a fungus. pieces and nodes are placed where the mycelia separate or merge (see Figure 9). The dataset contains several fungus species. Fungi are grown in various conditions and are exposed to attacks of grazers of various types. Each type of grazer modifies the growth of the fungus, resulting in a change of the structural properties of the network [boddy2010fungal] (see the structural changes between the fungi grown with or without grazers in Figure 9). For more consistency through the dataset and for computational time considerations, we select a subset of the graphs available in fricker2025fungi, by keeping graphs with a number of nodes smaller than 1500, resulting in a data set of 128 graphs, spanning three fungus species (Phallus impudicus, Phanerochaete velutina, Resinicium bicolour). The nodes of the graph are given with 2D spatial positions. The edges are weighted with biology-inspired resistance values that indicate how easy it is to transport nutrient along the given edge. This resistance value is proportional to the length of the edge and inversely proportional to the cross-section area.
@@
We apply our topological spatial graph coarsening algorithm to each graph of the dataset. Note that here, as the graph is given with specific weight that are relevant for the biological modeling, the coarsening is perform by using the resistance weights instance of edge lengths. Observe that the resistance is still a spatially and geometrically inspired quantity, similar to the length. Thus it is relevant to use it in this biological context. For each graph $G=(V,E,X)$ in the dataset, we apply our topological spatial graph coarsening procedure but replacing the use of the edge lengths the specific resistance values. Therefore, the scoring function is optimized over a grid $\Theta_{m}=\{q_{\alpha}(\mathcal{W}),\alpha\in\{1/(m+1),\dots,m/(m+1)\}\}$ adapted to the values of the resistance weights $\mathcal{W}=\{w_{u,v},\ (u,v)\in E\}$. In the experiments, $m$ is set to 10. Once we obtain the minimizer $\theta^{*}$ of $S$ over $\Theta_{m}$, we can save the corresponding quantile level $\alpha^{*}$ such that $q_{\alpha^{*}}(\mathcal{W})=\theta^{*}$. This $\alpha^{*}$ corresponds to the fraction of edges that are collapsed when the coarsening is performed with threshold $\theta^{*}$. The histogram of these optimal quantile levels $\alpha^{*}$ over the dataset are given in Figure 11. We can observe that based on the topology of the graph, our coarsening algorithm chooses to reduce the graphs by different amounts. Nonetheless, it is able to significantly reduce the graph sizes, as indicated by the boxplots representing the graph size distribution on both the original graph dataset and the dataset of reduced graphs in Figure 11. Examples of the obtained coarsened graphs and the corresponding score functions are shown from Figure 13 to 17.
@@
Figure 12: Illustration of the topological spatial reduction on a Phallus impudicus fungus without grazing.
@@
Figure 14: Illustration of the topological spatial reduction on a Phallus impudicus fungus with grazing from collembola (small grazers).
@@
Figure 16: Illustration of the topological spatial reduction on a Phallus impudicus fungus with grazing from millipedes (large grazers).
@@
To evaluate how well the information has been preserved in the reduced graph, we conduct the following experiment. We perform a classification task to predict the type of grazers that has attacked the fungus on both the original and the coarsened dataset. We then compare the accuracy of both classifiers to estimate how much of the original information is retained. For this experiment, we consider 3 types of grazing scenarios (3 classes): no grazer (control group), small grazers (< 10mm) including collembola, mites and nematodes, and large grazers (> 10mm) including millipedes, woodlices, and enchytraeid worms. To perform the classification, for each spatial graph we extract simple features from their persistent diagrams: number of connected components, mean, maximal and total persistence (where the persistence of a point in the diagram is the difference between its death and birth value) for the 1D persistence diagram, mean birth and death values for the 1D persistence diagram, $L_{2}$ norm of the concatenated 5 first persistence landscapes [bubenik2015statistical], and number of degree 1 nodes in the graph. For each dataset (original and reduced), we extract these features. A random forest classifier is trained for each dataset, with 200 trees. Performances are evaluated by the accuracy on a 10-fold cross validation. The accuracies of the classifiers are reported in Figure 18. The error bars represent the 95% confidence interval obtained by bootstrapping $10^{4}$ times the predictions on each sample given by the cross-validation procedure. Here, one should note that the dataset considered is not very large, therefore performances of the random forest is not expected to be very high. Nonetheless, we see that our learning procedure is able to capture relevant information as the obtained accuracies are significantly above 1/3, which would correspond to a totally uninformative random classification. Secondly, and more importantly, we observe only a slight accuracy drop between the original and reduced graphs, that does not seem to be statistically significant, as indicated by the overlapping confidence intervals. This clearly indicates that our coarsening algorithm has been able to significantly reduce the graph sizes by roughly a factor 2 (see Figure 11) while still preserving the key information contained in these spatial graphs such that learning algorithms can still perform as good as when applied on the original data. We emphasize here that we did not tune the amount of reduction applied by our coarsening method to preserve classification performance as the whole point of our proposed approach is to have a self-tuning coarsening algorithm. The reduction level is set by preserving the topology of the graph, which is often carrying meaningful information, as demonstrated in the present experiment.
@@ 7 Conclusion and further developments
We propose a new method for coarsening graphs with spatial coordinated associated to the nodes. The methods aims at preserving the topological properties of the graph. To capture the topological property of the spatial graph we adapt standard persistence diagram. We tested the method on few simulated data and one real world dataset of fungi. To grasp the trade off between retaining salient information and simplifying the graphs we perform a classification task on the fungi dataset. While the coarsening procedure significantly reduced the graph sizes, we observe only a small decrease in the accuracy, indicating that our coarsening approach has been able to preserve the key topological information used for the classification task.

### v2

@@ 6.3 Application on a classification task
We evaluate our topological and optimized coarsening procedure on a fungi network dataset. We consider a real-world dataset of fungal mycelia extracted from [fricker2025fungi]. Each fungus is represented as a spatial graph where the edges represent mycelia  [^1] The mycelium is the root-like structure of a fungus. branches and the nodes are placed where the mycelia branches separate or merge (see Figure 10). The dataset contains fungi of different species, grown in various conditions, and exposed to attacks of grazers of various types. Each type of grazer modifies the growth of the fungus, resulting in a change of the structural properties of the network [boddy2010fungal] For example, the structural changes between two fungi grown with or without grazers are shown in Figure 10). For more consistency through the dataset and for computational time considerations, we select a subset of the graphs with a number of nodes smaller than 1500, resulting in a data set of 128 graphs, spanning three fungus species (Phallus impudicus, Phanerochaete velutina, Resinicium bicolour). The nodes of the graph are given with 2D spatial positions and the edges are weighted with biological-inspired resistance values that indicate how easy it is to transport nutrient along the given edge. This resistance value is proportional to the length of the edge and inversely proportional to the cross-section area.
@@
We apply our topological spatial graph coarsening algorithm to each graph of the dataset. Note that here, as the graph is given with specific weight that are relevant for the biological modeling, the coarsening is perform by using the resistance weights instance of edge lengths. Observe that the resistance is still a spatially and geometrically inspired quantity, similar to the length. Thus it is relevant to use it in this biological context. For each graph $G=(V,E,X)$ in the dataset, we apply our topological spatial graph coarsening procedure but replacing the use of the edge lengths the specific resistance values. Therefore, the scoring function is optimized over a grid $\Theta_{m}=\{q_{\alpha}(\mathcal{W}),\alpha\in\{1/(m+1),\dots,m/(m+1)\}\}$ adapted to the values of the resistance weights $\mathcal{W}=\{w_{u,v},\ (u,v)\in E\}$. In the experiments, $m$ is set to 10. Once we obtain the minimizer $\theta^{*}$ of $S$ over $\Theta_{m}$, we can save the corresponding quantile level $\alpha^{*}$ such that $q_{\alpha^{*}}(\mathcal{W})=\theta^{*}$. This $\alpha^{*}$ corresponds to the fraction of edges that are collapsed when the coarsening is performed with threshold $\theta^{*}$. The histogram of these optimal quantile levels $\alpha^{*}$ over the dataset are given in Figure 12. We can observe that based on the topology of the graph, our coarsening algorithm chooses to reduce the graphs by different amounts. Nonetheless, it is able to significantly reduce the graph sizes, as indicated by the boxplots representing the graph size distribution on both the original graph dataset and the dataset of reduced graphs in Figure 12. Examples of the obtained coarsened graphs are shown in Figures 14, 16 and 18, with the corresponding score functions in Figures 14, 16, and 18.
@@
Figure 13: Illustration of the topological spatial reduction on a “Phallus impudicus“ fungus without grazing.
@@
Figure 15: Illustration of the topological spatial reduction on a “Phallus impudicus“ fungus with grazing from “collembola (small grazers).
@@
Figure 17: Illustration of the topological spatial reduction on a “Phallus impudicus“ fungus with grazing from millipedes (large grazers).
@@
To evaluate how well the information has been preserved in the reduced graph, we conduct the following experiment. We perform a classification task to predict the type of grazers that has attacked the fungus on both the original and the coarsened dataset. We then compare the accuracy of both classifiers, to estimate how much of the principal original information the coarsening method was able to retain. The experiment is based on the fact that the glazer type effect the grow of the fungal mycelium network [fricker2025fungi]. For this experiment, we consider 3 types of grazing scenarios (3 classes): no grazer (control group), small grazers (< 10mm) including collembola, mites and nematodes, and large grazers (> 10mm) including millipedes, woodlices, and enchytraeid worms. To perform the classification, for each spatial graph we extract simple features from their persistent diagrams: number of connected components, mean, maximal and total persistence (where the persistence of a point in the diagram is the difference between its death and birth value) for the 1D persistence diagram, mean birth and death values for the 1D persistence diagram, $L_{2}$ norm of the concatenated 5 first persistence landscapes [bubenik2015statistical], and number of degree 1 nodes in the graph. For each dataset (original and reduced), we extract these features. A random forest classifier is trained for each dataset, with 200 trees. Performances are evaluated by the accuracy on a 10-fold cross validation. The accuracies of the classifiers are reported in Figure 19. The error bars represent the 95% confidence interval obtained by bootstrapping $10^{4}$ times the predictions on each sample given by the cross-validation procedure. Here, one should note that the dataset considered is not very large, therefore performances of the random forest is not expected to be very high. Nonetheless, we see that our learning procedure is able to capture relevant information as the obtained accuracies are significantly above 1/3, which would correspond to a totally uninformative random classification. Secondly, and more importantly, we observe only a slight accuracy drop between the original and reduced graphs, that does not seem to be statistically significant, as indicated by the overlapping confidence intervals. This clearly indicates that our coarsening algorithm has been able to significantly reduce the graph sizes by roughly a factor 2 (see Figure 12) while still preserving the key information contained in these spatial graphs such that learning algorithms can still perform as good as when applied on the original data. We emphasize here that we did not tune the amount of reduction applied by our coarsening method to preserve classification performance as the whole point of our proposed approach is to have a self-tuning coarsening algorithm. The reduction level is set by preserving the topology of the graph, which is often carrying meaningful information, as demonstrated in the present experiment.
@@ 7 Conclusion and further developments
We propose a new method for coarsening graphs with spatial coordinated associated to the nodes. The methods aims at preserving the topological properties of the graph, captured by adapting standard persistence diagram. We prove the novel coarsening method is equivariant with respect to the action of the symmetric group. We tested the method on few simulated data and two real world datasets: a road network and a set of fungal mycelia. To grasp the trade off between retaining salient information and simplifying the graphs we perform a classification task on the fungi dataset. While the coarsening procedure significantly reduced the graph sizes, we observe only a small decrease in the accuracy, indicating that our coarsening approach has been able to preserve the key topological information used for the classification task.
@@
Acknowledgments
We would like to thank Mark Fricker for the valuable discussions on the fungus data and for providing us with the dataset. We also want to acknowledge the financial support of the CNRS-Imperial "Abraham de Moivre" International Research Laboratory on this project.
Statements and Declarations
Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.
Funding Partial financial support was received from the CNRS-Imperial "Abraham de Moivre" International Research Laboratory Travel fund.